# Generative Active Adaptation for Drifting and Imbalanced Network Intrusion Detection

### v1

@@ 4.3.2. UGR’16 Datasets for July and August
The performance metrics in Table 6 confirm the presence of concept drift in the UGR’16 dataset, as the No Adapt (0%) setting causes the F1-score to drop to 68.77% (MLP) and 74.33% (XGBoost) when training on July data and testing on August data. With just 1% labeled data, NetGuard AL + Augmentation achieves 89.66% F1 (MLP) and 99.29% F1 (XGBoost), outperforming all baselines and nearing Full Adaptation (93.43% MLP, 99.55% XGBoost), proving its efficiency in mitigating drift. The class-wise performance results for Spam and Botnet with data augmentation are provided in the Appendix.
@@
Table 5. Comparison with OOD sample selection using CADE
Method F1 (%) FPR FNR Label Cost
No Adapt 60.62 0.5495 0.3062 0%
CADE + Retrain 72.26 0.2748 0.1531 40%
NetGuard (End-to-end) 81.78 0.2686 0.1419 1%
Table 6. NetGuard ’s end-to-end performance on UGR’16 (Train: July $\rightarrow$ Test: August)
MLP XGBoost
Method Label F1 (%) FPR FNR F1 (%) FPR FNR
No Adapt 0% 68.77 0.08 0.3 74.33 0.062 0.246
Uncertainty 0.1% 67.68 0.0755 0.2923 79.40 0.0466 0.1855
0.5% 69.02 0.0718 0.2790 85.73 0.0353 0.1403
1% 68.40 0.0735 0.2864 90.36 0.0253 0.1003
CLUE 0.1% 15.75 0.1981 0.7901 89.04 0.0266 0.0800
0.5% 13.82 0.2002 0.8005 92.31 0.0188 0.0561
1% 36.91 0.1474 0.5719 94.04 0.0149 0.0446
NetGuard AL (Ours) 0.1% 67.43 0.0772 0.3006 79.54 0.0462 0.1839
0.5% 67.97 0.0752 0.2929 85.24 0.0362 0.1439
1% 70.57 0.0670 0.2615 90.64 0.0247 0.0976
NetGuard (End-to-end) 1% 89.66 0.0261 0.0822 99.29 0.0019 0.0065
Full Adapt 100% 93.43 0.0167 0.05 99.55 0.0012 0.0038
4.4. Ablation
We conduct an ablation study by evaluating the impact of different generative models (RTF vs. TVAE) on NetGuard end-to-end classification. We evaluate this on the UGR16 dataset with 1% sampling budget and MLP classifier. We also analyze effects of different data generation amounts (3× vs. 10× augmentation) on Spam and Botnet classes (illustrated in Appendix).
Effects of different generative models.
The left graph in Fig. 3 shows that RTF improves F1-score more than TVAE, with 75.33% vs. 40.66% (Botnet) and 96.26% vs. 94.01% (Spam), indicating better augmentation effectiveness. The right graph shows that RTF has a lower W2 distance, meaning its synthetic samples are closer to the real distribution (0.0059 vs. 0.0133 for Botnet, 0.0340 vs. 0.2292 for Spam). This confirms that RTF generates higher quality data, leading to better classification performance.
5. Engineering Lessons
Lesson-1: Hyper-parameters of active learning and data augmentation.
The end-to-end effectiveness of NetGuard relies on optimizing the labeling budget for active learning, the volume of synthetic data for augmentation, and the types of classes selected for augmentation. As observed in Table 3 and Table 6, choosing a 1% labeled data budget achieved near full adaptation performance, highlighting the need for efficient annotation strategies. For the other tunable knobs of NetGuard ’s augmentation module, we select classes for augmentation that continue to exhibit poor performance even after NetGuard ’s active learning stage, where this under-performance is primarily due to scarcity of available data for that class. While increasing augmentation (e.g., 3× vs. 10×) enhances intrusion detection, an excessive augmentation may lead to redundancy and model overfitting causing an increase in the false positive rate. Therefore, an optimal augmentation level should be empirically determined based on dataset characteristics and minority class distributions.
Lesson-2: Demonstrating adaptability to temporal and spatial concept drift.
NetGuard effectively handles concept drift across both spatial and temporal dimensions. For spatial drift, we evaluated CIC-IDS 2017 (small network) vs. CIC-IDS 2018 (large network), simulating real-world network evolution, where traffic patterns shift as infrastructure scales. NetGuard improved end-to-end F1-score from 0.60 to 0.86, proving adaptability across architectures. For temporal drift, UGR’16 (real-world ISP dataset) captured periodic traffic variations over months, improving F1-score from 0.69 to 0.89. We observed consistent performance even when training in July and testing in later months, highlighting NetGuard ’s robustness in evolving networks, ensuring resilience to both expanding infrastructures and time-dependent traffic shifts.
Lesson-3: Impact of downstream classifiers.
We evaluate NetGuard across two categories of classifiers—MLP (deep learning) and XGBoost (gradient boosting)—confirming that both are affected by concept drift but improve with better data selection. Critically, NetGuard does not require any modifications to the classifier itself; instead, it enhances performance by mitigating concept drift with data scarcity issues through active learning and generative augmentation. For UGR’16, both MLP and XGBoost showed improvements, though MLP benefited more due to its reliance on training data, while XGBoost, being an ensemble model, is inherently more robust to class imbalance. This demonstrates that targeted data selection strategies can significantly enhance classifier performance, regardless of model type.
6. Discussion
We now discuss some actionable points on NetGuard including our assumptions, existing limitations, and future prospects.
Generalization to other domains in cybersecurity.
The NetGuard framework is designed to handle concept drift and imbalanced data, making it applicable beyond network intrusion detection. It can be extended to malware classification, fraud detection, smart grid attack detection, and IoT anomaly detection, where data distributions evolve over time. For instance, in malware detection, NetGuard ’s density-aware active learning can help adapt models to new malware variants, while generative augmentation can synthesize rare attack patterns.
Correct observation labels.
We assume that the labels for benign and attack traffic in our datasets are accurate, but real-world datasets often contain label noise due to manual annotation errors or misclassification by automated systems. Prior research has reported labeling inconsistencies in CIC-IDS 2017/2018, primarily due to delayed attack detection (Liu et al., 2022). However, since both datasets apply the same detection strategies in a controlled manner, these inconsistencies do not affect our study, as the labeling methodology remains consistent across both datasets. In the future, we aim to explore semi-supervised learning methods to reduce reliance on fully labeled datasets and incorporate reviewer feedback technique with active learning for label correction. For example, the model can prioritize samples with uncertain or conflicting labels for review, allowing the reviewer to correct any errors (Kremer et al., 2018).
Adversarial machine learning.
As NetGuard relies on active learning and generative models, it could be targeted by adversarial attacks aimed at misleading sample selection or poisoning data augmentation (Biggio and Roli, 2018). Attackers might manipulate traffic to avoid detection by introducing subtle variations closely resemble the previous benign traffic flows. Future work could integrate adversarial training techniques, the model could be made more robust to such attacks. Additionally, the generative models used for data augmentation could be leveraged to generate adversarial examples, which could then be used to train the model to be more resilient.
@@ 7. Conclusion
We proposed a generative adaptation framework NetGuard that combines distribution-aware active learning with deep generative model-based augmentation to address concept drift in NIDS. Our approach improves model adaptability to evolving threats, particularly for rare attack classes like Botnet. By leveraging active learning, we enhance sample selection efficiency, while deep generative models contribute to data diversity, mitigating the effects of data imbalance. Evaluation on CIC and UGR’16 NIDS datasets demonstrated significant performance gains, highlighting the effectiveness of our method in dynamic network environments. The integration of these techniques enables efficient, scalable, and adaptive NIDS, reducing reliance on costly full adaptation while improving detection of both common and rare attacks. Beyond adaptation, our framework has the potential to enhance the robustness of both training and evaluation processes in NIDS. This generative active approach can be leveraged to create diverse and realistic adversarial scenarios.improving model generalization to unseen attacks, to further improve ML-based NIDS in adversarial environments.
@@ Appendix A Appendix
Effects of data generation amount
@@
Overall adaptive effectiveness in CIC Train: 2018 → Test: 2017 setting:
In the reverse case for CIC datasets shown in Table 8, augmentation was applied only to Infiltration, as it had the lowest classification performance after the NetGuard AL stage. The results show that Infiltration’s F1-score improved from 0.02 with NetGuard AL to 0.39 with augmentation, nearly matching Full Adapt (0.40). This overall validates that NetGuard AL combined with augmentation reduces the dependency on fully labeled datasets by bringing model performance closer to the Full Adaptation scenario.
@@
Table 7. Class-wise F1-score of NetGuard usingMLP and XGBoost classifiers on UGR’16 (Train: July $\rightarrow$ Test: August)
@@
Table 8. Class-level F1-score for Train: 2018$\rightarrow$ Test: 2017
Attack No Adapt 1% CLUE 1% NetGuard AL End-to-end Full Adapt
(NetGuard AL+Data Aug.)
@@
Figure 5. Comparison of latent space visualizations for different active learning sampling strategies

### v2

@@ 4.3.3. UGR16:
Overall adaptation effectiveness (Train: July $\rightarrow$ Test: August). The performance metrics in Table 5 confirm the presence of concept drift in the UGR’16 dataset, as the No Adapt (0%) setting causes the F1-score to drop to 68.77% (MLP) and 74.33% (XGBoost) when training on July data and testing on August data. With just 1% labeled data, NetGuard PS + Augmentation achieves 89.66% F1 (MLP) and 99.29% F1 (XGBoost), outperforming all baselines and nearing Full Adaptation (93.43% MLP, 99.55% XGBoost), proving its efficiency in mitigating drift.
Table 5. Performance on UGR’16 (Train: July $\rightarrow$ Test: August)
MLP XGBoost
Method Label F1 (%) FNR FPR F1 (%) FNR FPR
No Adapt 0% 68.77 0.300 0.08 74.33 0.246 0.062
Uncertainty 0.1% 67.68 0.2923 0.0755 79.40 0.1855 0.0466
0.5% 69.02 0.2790 0.0718 85.73 0.1403 0.0353
1% 68.40 0.2864 0.0735 90.36 0.1003 0.0253
CLUE 0.1% 15.75 0.7901 0.1981 89.04 0.0800 0.0266
0.5% 13.82 0.8005 0.2002 92.31 0.0561 0.0188
1% 36.91 0.5719 0.1474 94.04 0.0446 0.0149
NetGuard 1% 89.66 0.0822 0.0261 99.29 0.0065 0.0019
Full Adapt 100% 93.43 0.0500 0.0167 99.55 0.0038 0.0012
4.4. Ablation Study
We evaluate generative model choices (RTF vs. TVAE) in NetGuard using UGR’16 (1% sampling budget, MLP classifier). Comparative analysis includes: (1) augmentation strength (3× vs. 10×) for Spam/ Botnet classes in UGR’16 (see Appendix 4), and (2) synthetic data quality via PCA visualizations (CIC-IDS, Appendix 5). We compare RTF and TVAE as generative backbones within NetGuard . As shown in Figure 3, the left graph depicts end-to-end classification performance, where RTF outperforms TVAE in both classification (F1: 75.33% vs. 40.66% for Botnet; 96.26% vs. 94.01% for Spam) and the right graph shows statistical fidelity (W2 distance as: 0.0059 vs. 0.0133 for Botnet; 0.0340 vs. 0.2292 for Spam). Our W2 scores surpass NetShare’s reported EMD values for UGR16 dataset ($\approx$0.10–0.15) (Yin et al., 2022). For CIC-IDS 2018, our normalized W2 distances between synthetic and original samples/class are 0.34 (Infiltration), 0.92 (Web Attack), 0.03 (FTP-BruteForce)—reflect class-specific fidelity, comparable to NetShare’s reported EMD values.
@@ 5. Conclusion
NetGuard is a novel generative framework for NIDS that combines distribution-shift aware sample selection with conditional data augmentation to address concept drift and class imbalance. Our method conditions synthetic sample generation on drift patterns to better adapt to evolving threats, particularly rare attacks like web attack and botnet. Evaluations on CIC and UGR’16 datasets demonstrate improved detection in dynamic environments while reducing adaptation costs. The framework’s generative approach maintains data diversity and enhances robustness against both common and emerging attacks.
@@ Appendix A Appendix
Effects of data generation amount on UGR’16
@@
Table 6. Class-wise performance of MLP and XGBoost for adaptation methods on UGR’16 (Train: July $\rightarrow$ Test: August)
@@
Class-specific adaption effectiveness in CIC Train: 2018 → Test: 2017 setting.
In the reverse case for CIC experiments as shown in Table 7, augmentation was applied only to Infiltration, as it had the lowest classification performance after the NetGuard PS stage. The results show that Infiltration’s F1-score improved from 0.02 with NetGuard PS to 0.39 with augmentation, nearly matching Full Adapt (0.40). This overall validates that NetGuard PS combined with augmentation reduces the dependency on fully labeled datasets by bringing model performance closer to the Full Adaptation scenario.
[…]
Attack No Adapt 1% CLUE 1% NetGuard PS End-to-end Full Adapt
(NetGuard PS+Data Aug.)
@@
Sensitivity analysis for GMM-based prior selection.
Table 8 illustrates that using 1% labeling, a GMM with K=10 and diagonal covariance achieves the best trade-off between F1-score and runtime across both CIC-IDS 2017/18 and UGR’16, making it the optimal setting for prior selection under drift; its parametric nature enables efficient, scalable modeling of distributional shifts, unlike non-parametric methods which are often computationally expensive and sensitive to hyperparameter tuning.
Table 8. Sensitivity of GMM parameters at 1% labeling
Dataset K Covariance F1 Runtime (s)
CIC-IDS 2017/18 5 diagonal 0.81 3.48
10 diagonal 0.84 5.97
10 full 0.83 12.96
UGR’16 (Train: July $\rightarrow$ Test: Aug.) 10 diagonal 0.71 6.32
Computational cost of NetGuard.
Table 9 presents a breakdown of the computational usage of various NetGuard components. NetGuard’s REaLTabFormer training takes $\approx$40 minutes per cycle on an A100 GPU (avg. 25% utilization), while sample generation costs 4–6 minutes per 1,000 samples (0.24–0.36s/sample, ¡5% GPU). GMM-based prior selection takes 6 seconds per 1,000 samples ($\approx$6ms/sample, CPU-bound).
Table 9. Computational Overhead of NetGuard
Component Cost / 1k Samples Latency / Sample Hardware Util.
GMM-based prior selection 5.97–6.32 sec $\approx$ 6 ms CPU-bound / ¡5%
RTF training (initialization)  40 min GPU /  25%
RTF sample generation 4–6 min 0.24–0.36 sec GPU / ¡5%
PCA visualization of attack classes before and after data augmentation
We provide PCA visualizations of CIC-IDS 2018 attack samples before (left) and after (right) 3× data augmentation in Fig.5. The augmented samples closely follow the original distribution while remaining distinguishable across classes.
@@
Figure 5. PCA projection of four CIC-IDS 2018 attack classes before (all samples, left) and after (3× augmentation on only the PS samples, right).
Evolving trustee rules on CIC 2017-¿ 2018 dataset
Table 10 shows the evolving Trustee classification rules for Infiltration, Web Attack, and FTP-BruteForce under different adaptation scenarios on CIC-IDS 2017→2018. Our analysis of classifier decision rules (Table 10) demonstrates that NetGuard ’s prior selection (1% labeling) achieves 89% feature commonality with full target domain adaptation (100%), validating its ability to identify the most critical drift indicators (e.g., Fwd IAT Mean thresholds, PSH flag counts) with minimal supervision. The detailed rule breakdown in the Appendix reveals consistent preservation of core detection logic across all attack classes - Infiltration maintains flow activity thresholds, Web Attacks adapt packet timing features while preserving structural patterns, and FTP-BruteForce retains segment size requirements - confirming that NetGuard ’s efficient approach captures nearly all discriminative features of expensive full adaptation. This strong alignment between limited and full adaptation rules proves NetGuard ’s effectiveness for practical deployment where exhaustive retraining is infeasible.
Table 10. Trustee classification rules (Train:2017 $\rightarrow$ Test:2018)
Class No Adaptation NetGuard PS (1%) Full Adaptation (100%)
Infiltration PSH Flag Count ¿ 0.16
Active Mean ¿ 20.76
Dest Port $\leq$ 0.30, Bwd Header Len $\leq$ 0.01, Flow IAT Mean $\leq$ 0.33, min_seg_size_fwd $\leq$ 0.00
Active Min ¿ 3.64, Fwd IAT Max $\leq$ 0.23
Web Attack Fwd IAT Mean ¿ 0.33, Fwd IAT Total $\leq$ 0.53, Fwd IAT Min $\leq$ 0.18
Flow Packets/s ¿ 0.27, Flow Duration $\leq$ 0.56
Init Win Bytes Fwd ¿ 0.72, Fwd IAT Mean $\leq$ 0.29, Flow Bytes/s $\leq$ 0.05
Fwd Packet Len Std ¿ 1.37
FTP-BruteForce Bwd Packets/s min_seg_size_fwd ¿ 0.00, Flow Duration $\leq$ 0.56, Packet Len Var $\leq$ 0.39
Flow IAT Max ¿ 0.49, Flow IAT Mean $\leq$ 0.33
min_seg_size_fwd ¿ 0.00
Fwd Packets/s ¿ 0.16
@@
Figure 6. Comparison of latent space visualizations for different prior selection sampling strategies